OpenStudy (anonymous):

PLEASE HELP!!! solution to the linearization around zero of the system x'=−2x−y−3x^5 y'=9x−2y−x(y^3) X= Y=

OpenStudy (athe):

For linear approximation around zero \[3x^5, x\cdot y^3\rightarrow 0\]: \[x'=-2x-y \\ y'=9x-2y\]